In 2019, Lamar, a top college quarterback, signed a $40.4 million football contract structured as follows: for the years 2019 through 2024 inclusive, it paid him equal annual upfront payments which totaled $3.9 million. He also received an upfront signing bonus of $500K. The balance of the contract was paid out equally for 30 years (end of year) following the initial series of annual payments.

On the other hand, Melvin, a star running back, signed a deal which paid him $18 million over four years as follows: 50% upfront with the balance of the contract paid out equally at the end of each year for the duration of the contract.

Around the same time J.J., a bone crushing lineman, received a 4-year, $24 million contract extension which included an upfront signing bonus of $6 million; $1.5 million/year in salary for the duration of the contract and a $300K payment at the beginning of each year for the next 40 years.

Assume annual interest compounding, a discount rate of 10% and no taxation impact.                                  

Who got offered the better deal and why?  
Answer #1

We can compare the deals by comparing their present value at 10%

For Lamar

upfront payment (year 0 ) = $0.5 million + $3.9/6 million = $1.15 million

Year (1-5) = $3.9/6 million = $0.65 million

For next 30 years (year 6- 35) , he got = $(40.4-0.5-3.9)/30 million each year =$1.2 million each year

So, Present value of his cash flows

= 1.15+ 0.65/1.1+...+0.65/1.15+ 1.2/1.16+....+1.2/1.135

=$10.638 million

For Melvin,

upfront payment (year 0 ) =50% of $18 million =$9 million

Year (1-4) = $9/4 million = $2.25 million

So, Present value of his cash flows

= 9+2.25/1.1+ ...+2.25/1.14

= $16.13 million

For J.J,

upfront payment (year 0 ) = $6million

Year (1-4) = $1.5 million

For next 40 years (year 5- 44) , he got $0.3 million each year

So, Present value of his cash flows

= 6+ 1.5/1.1+...+1.5/1.14+ 0.3/1.15+....+0.3/1.144

=$12.76 million

From the above, we can say that Melvin got the best deal as he is getting the highest amount in terms of present value.